OpenStudy (anonymous):

how do sove for the solutions for ; the square root of x+ the square root of x-36 which then =0?

OpenStudy (anonymous):

Subtract the square root of x-36 from both sides! now: the square root of x = -the square root of x-36

OpenStudy (anonymous):

\[\sqrt{x} = -\sqrt{x-36}\]

OpenStudy (anonymous):

what about the 2?

OpenStudy (anonymous):

Square both of them\[(\sqrt{x})^{2} = (-\sqrt{x-36})^{2}\]

OpenStudy (anonymous):

what 2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im so sry the whole thing is equal to 2 not 0 im so sry

OpenStudy (anonymous):

ok...then it is: \[(\sqrt{x})^{2} = (2-\sqrt{x-36})^{2}\]

OpenStudy (anonymous):

\[x = 4 - 2\sqrt{x-36} - 2\sqrt{x-36} + x - 36\]

OpenStudy (anonymous):

\[x = -32 + x + -4\sqrt{x-36}\]

OpenStudy (anonymous):

\[x - x + 32 = -4\sqrt{x-36}\]

OpenStudy (anonymous):

\[32 = -4\sqrt{x-36}\]

OpenStudy (anonymous):

\[(32)^{2} = (-4\sqrt{x-36})^{2}\]

OpenStudy (anonymous):

no no!!

OpenStudy (anonymous):

wait a sec!

OpenStudy (anonymous):

Instead of square rooting 32 and the other, first divide -4 from both sides!

OpenStudy (anonymous):

\[(-8)^{2} = (\sqrt{x-36})^{2}\]

OpenStudy (anonymous):

you following?

OpenStudy (anonymous):

64 = X-36 64 + 36 = X - 36 + 36 100 = X

OpenStudy (anonymous):

thank you! :)